# High Performance Distributed Control Law for Large-Scale Linear Systems: A Partitioned Distributed Observer Approach

### v3

@@ Abstract
In recent years, the distributed-observer-based distributed control law has shown powerful ability to arbitrarily approximate the centralized control performance. However, the traditional distributed observer requires each local observer to reconstruct the state information of the whole system, which is unrealistic for large-scale scenarios. To fill this gap, this paper develops a greedy-idea-based large-scale system partition algorithm, which can significantly reduce the dimension of local observers. Then, the partitioned distributed observer for large-scale systems is proposed to overcome the problem that the system dynamics are difficult to estimate due to the coupling between partitions. Furthermore, the two-layer Lyapunov analysis method is adopted and the dynamic transformation lemma of compact errors is proved, which solves the problem of analyzing stability of the error dynamic of the partitioned distributed observer. Finally, it is proved that the distributed control law based on the partitioned distributed observer can also arbitrarily approximate the control performance of the centralized control law, and the dimension of the local observer is greatly reduced compared with the traditional method. The simulation results show that when the similarity between the physical network and the communication network is about $80\%$, the local observer dimension is greatly reduced by $90\%$ and the relative error between the performance of the distributed control law and that of the centralized control law is less than $1\%$.
@@ I Introduction
Distributed observer is a cooperative observer network composed of $N$ agents and a communication network. Each agent is a node of the network and contains a local observer as well as a local controller. Each local observer only has access to partial system outputs, and all local observers cooperate to complete the state estimation of the whole system through information interaction via communication networks. Distributed observer has extensive application value and profound theoretical value. In terms of application value, global scholars have enabled distributed observer to play important roles in flexible structures [1], smart vehicles [2], microgrids [3, 4], deep-sea detectors [5, 6], spacecraft in low Earth orbit [7] and other fields. In terms of theoretical value, distributed observer can help distributed control law to improve its performance with a qualitative change. Readers may refer the theoretical value to the article of [2, 8, 9], [10], and [11], which prove that the distributed-observer-based distributed control law has the same performance as that of the centralized control law. Owing to its broad application prospects and outstanding theoretical significance, distributed observer finds itself a center of research in recent years [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 24, 25, 26].
It should be noted that the current distributed observer has an obvious shortage: It requires that every agent has to reconstruct the states of the whole system. In other words, if the system dimension is $n$, then the dimensions of local observers on every agent are all $n$, which is the origin of a special term in the field of distributed observer: state omniscience. Accordingly, it is barely able for a large-scale system to achieve state omniscience because of the high dimensions of local observers. After all, we cannot require each agent to compute a local observer with the same dimension as the large-scale system, otherwise, we will lose the original intention of the distributed design. Although a few articles have studied the minimum order distributed observer [28] or reduced order distributed observer [29], they replace $N$ full-dimension observers contained in the local observer with $N$ reduced-dimension observers. It cannot essentially solve the problem of the high dimension of local observers caused by large-scale systems.
This paper will explore whether there is a method that can effectively reduce the dimension of local observers on each agent. One of the biggest challenges in studying this problem is to reduce the dimension of local observers without affecting their ability to help distributed control approach centralized control performance. According to [30], we know that the reason why distributed control law cannot achieve the same performance as that of the centralized control law is that distributed control law cannot obtain as much state information as centralized control. However, in [2], an agent can estimate the whole system states through the state omniscience of distributed observer. It indicates that state omniscience is an important factor in why the distributed-observer-based distributed control law has centralized control performance. Therefore, approaching the centralized control performance without relying on state omniscience presents the distributed-observer-based distributed control law with huge challenges.
The idea to overcome these challenges in this paper comes from the thinking of large-scale interconnected systems structure. We find that, for large-scale interconnected systems (it is better to use a graph to represent the interconnection relationship among subsystems and name this graph as the physical network), the local control law on each agent does not actually need the whole large-scale system states. Generally speaking, the states of physical network neighbor agents are enough for an agent to achieve centralized control performance. For example, the states employed in the centralized control law of interconnected microgrid systems are only the states of each agent’s physical network neighbor agents. Inspired by this, can we build a novel distributed observer so that the local observer on each agent only needs to estimate the states of its physical neighbor agents? If feasible, the local observer dimension on each agent can be greatly reduced because the interconnection of a large system is often sparse, and the average number of physical neighbor agents of each agent is generally far less than the number of all agents. In addition, since the states of all physical neighbor agents of an agent are estimated, the local control law can use the same state information as that of centralized control. It indicates that the aforementioned method does not affect the performance of the distributed observer and its role in distributed control law while reducing the dimension of all local observers.
Building the above-mentioned new distributed observer is a formidable task because a local observer containing only the physical network neighbor states can be established if and only if the physical network and communication network are identical. Unfortunately, physical networks and communication networks are generally different. For this reason, this paper tries to develop a method to partition all nodes. Each partition contains multiple nodes, and each node belongs to multiple partitions. Then, distributed observer can be designed in each partition, and all agents can complete the state estimation of all their physical neighbors through one or more partitioned distributed observers. Hence, the main focus of the partition is on enabling the union of all partitions on an agent to cover all its neighbor agents (this is a necessary condition to ensure that the distributed-observer-based distributed control law has centralized control performance), and also requiring the number of agents in the union of partitions to be as small as possible (reduce the dimension of local observers on each agent as much as possible).
Therefore, the first scientific challenge faced by this paper is how to establish an intelligent partition method that meets the above conditions. There is a lot of research on large-scale network partition in the existing literature, but these methods are based on one or more of the following conditions: 1) No intersection between any two partitions [31, 32, 33]; 2) One needs to know how many partitions in advance [34]; 3) There is a clear and single objective function [35, 33, 36]; 4) Partition the node coverage area [37, 38]. However, the partition problem in this paper does not meet any of the above conditions. Furthermore, the existing partition methods are all based on the topology information of only one network, while the topology information of two networks (physical network and communication network) should be considered at the same time in this paper. Therefore, the existing partition methods are not applicable to this work.
The second scientific challenge is how to deal with the problem of error dynamic coupling between various partitions. Although the idea of this paper is to partition large-scale systems and build distributed observer for small-scale systems within the partitions, distributed observers in the partitions are not independent, and they are dynamically coupled with the distributed observer in other partitions. If we do not eliminate this coupling, the observer is not implementable because they include the unknown states from other partitions. On the contrary, if this coupling is eliminated, we will face the model mismatch problem. What is more difficult is that even if the coupling relationship is eliminated in the observer design, it will occur in the error dynamics and lead to the difficulty of stability analysis. Therefore, the classical distributed observer theory, which is close to maturity, cannot be used either in the partitioned distributed observer design or in the error dynamic stability analysis.
Focusing on the above challenges, this paper contributes to the following three aspects:
1) We propose a reasonable and feasible partition method for large-scale systems. Furthermore, the partition method proposed in this paper can achieve the Pareto optimal solution (Each agent has its own most expected partitioning result, so the partitioning problem in this paper is a multi-objective optimization problem).
2) A design method is proposed for the partitioned distributed observer and a two-layer Lyapunov method is developed to analyze the stability of the error dynamic of the partitioned distributed observer. The proposed method guarantees the implementation of the partitioned distributed observer. It is also proved that each agent can estimate the states of all its physical network neighbor agents, and the estimation error can dynamically converge to any small invariant set.
3) Although the dimension of each local observer is greatly reduced, the distributed control law based on the partitioned distributed observer can still arbitrarily approximate the control performance of centralized control. It provides an important theoretical basis for the application of distributed observer in large-scale systems.
@@
The remainder of this paper is organized as follows. Section II formulates the problem. Section III gives the partition method and partitioned distributed observer design method. The performance of the error dynamics of partitioned distributed observer and closed-loop system are analyzed in Section IV. The simulation results are given in Section V, and Section VI concludes this paper.
[…]
@@
$\mathcal{O}_{p}$ The $p$th partition
$\mathscr{P}_{i}$ Set of partitions that containing node $i$
@@
$D(\mathscr{P}_{i})$ $\sum_{\mathcal{O}_{j}\in\mathscr{P}_{i}}\mathcal{O}_{j}$
$\hat{x}_{li}^{(p)}$ Estimation of $x_{i}$ generated by observers on agent $l$ in partition $\mathcal{O}_{p}$
@@ II-A Overall objectives
In traditional distributed observer, each local observer—the orange block in Figure 1—is required to reconstruct all the states of the entire system based on $y_{i}$ and the information exchanged via communication network. Then, each local controller—blue block in Figure 1—can use all the information of the entire system. Therefore, all local observers can jointly establish a distributed control law with centralized performance. Many existing studies, such as [2, 9], have proved this result, i.e., the distributed-observer-based distributed control law can achieve centralized control performance. However, in these studies, each local observer needs to reconstruct the states of the whole system, i.e, dimension of observers on each agent is $nN$. This is not realistic for large-scale systems. Therefore, this paper will investigate whether it is possible to achieve distributed control laws with arbitrary approximation of centralized control performance without requiring each local observer to estimate global information.
@@
1) Design the partitioned distributed observer with $\hat{x}_{i\star}$ being the states of the $i$th local observer, and $\hat{x}_{i\star}$ satisfies $dim\{\hat{x}_{i\star}\}\ll dim\{x\}$;
@@ II-B An illustrative example for the target problem
Figure 2: Partition examples of simple network systems.
@@
We take node $13$ as an example to demonstrate the specific issues to be studied in this paper. As seen in Figure 2, node $13$ has $5$ physical neighbors ($11,12,14,15,16$) and $3$ communication neighbors ($11,14,16$). To design a local control law to make the control performance as close as possible to the centralized control law, agent $13$ needs the states of all its physical neighbors. In traditional method, a distributed observer will be designed such that agent $13$ can reconstruct all states of the entire system. However, only the states of $11,12,14,15,16$ are required. Therefore, this paper intends to design a group of partitions and design distributed observer in each partition. In this way, the useless information estimated by each agent can be greatly reduced. For example, in Figure 2, node $13$ belongs to $3$ partitions—$\{11,12,13\}$, $\{13,15,16\}$, and $\{13,14\}$. By designing distributed observers separately in three partitions, agent $13$ only needs to estimate the states of $5$ subsystems (in traditional methods, agent $13$ needs to estimate the states of all $47$ subsystems). With partitioned distributed observers, local control laws can be established based on the results of state estimation.
Up to now, we have briefly described the main idea of this paper through an example. To achieve this idea, there are mainly three steps involved: 1) Design a reasonable and effective partitioning algorithm; 2) Design partitioned distributed observer in each partition; 3) Design distributed control law.
@@
The reader should note that Assumption 1 is not completely the same as the traditional assumption in distributed observers. In existing literature [15, 9], it is generally assumed that $(C,A)$ is observable, but $(C_{i},A)$ may not be observable. Assumption 1 used in this paper is given owing to the form of system (1)–(2). This system format is easier to express our core views on partitioning and status information retrieval. Actually, the partitioned distributed observer can also be achieved based on the General system $\dot{x}=Ax,~{}y=Cx$ and the traditional assumption as in [15, 9]. However, due to limitations in space and the need to express the main idea, the theory of partitioned distributed observers for general systems will be included in our future research.
III Design of partitioned distributed observer
@@
A large-scale network partition algorithm will be given in this section. Then, based on the partition results, we propose the design method and observer structure of the partitioned distributed observer.
III-A Network partition
@@
As mentioned in Section I and Subsection II-B, it is unnecessary to require each local observer to reconstruct the states of the whole large-scale system. Theoretically, each agent only needs to estimate part of states required by its local control law. Therefore, in order to establish a distributed observer that can only estimate the state information required by the local control law, we need to partition the network and ensure that all partitions of each agent contain all its physical network neighbors. To illustrate it in more detail, we introduce several simple examples.
Figure 3: Partition examples of simple network systems.
@@
In Figure 3a), there is a simple network system with $6$ nodes. One may recognize its partition by observation easily: it could be partitioned into three categories with $\{1,2,3\},\{4,5,6\}$, and $\{3,4\}$. With this partition, the physical network neighbors of all nodes have been included in their partitions. For example, $\mathcal{N}_{3}=\{1,2,4\}$, and partitions to which node $3$ belonging are $\{1,2,3\}$ and $\{3,4\}$. Obviously, $\mathcal{N}_{3}\subset\{1,2,3\}\cup\{3,4\}$. Similarly, Figure 3b) is a star network system with $9$ nodes. One can partition it with $\{1,2\}$, $\{1,3\}$, $\{1,4\}$, $\{1,5\}$, $\{1,6\}$, $\{1,7\}$, $\{1,8\}$, $\{1,9\}$ and all agents’ physical neighbors are covered by their partitions. As for Figure 3c), the acquisition of partitions is not so intuitive, but we can still partition it with $\{1,2,3,4\}$, $\{5,6,7,8\}$, and $\{3,5\}$ after the observation and analysis. Through the above examples, we know the basic idea and function of partition, but we cannot obtain reasonable partition results through observation for large-scale systems. Therefore, a reasonable partition algorithm is necessary for distributed observers of large-scale systems.
@@
To this end, we propose a partition method based on the greedy algorithm. Before showing the algorithm details, some notations are introduced. $\mathcal{O}_{k}\subset\mathcal{V}$ is defined as the $k$th partition of the large-scale system, and $\mathscr{P}_{i}$ denotes the set of partitions that containing node $i$. For example, in Figure 3c), $\mathcal{O}_{1}=\{1,2,3,4\}$, $\mathcal{O}_{2}=\{5,6,7,8\}$, and $\mathcal{O}_{3}=\{3,5\}$. In addition, $\mathscr{P}_{1}=\{\mathcal{O}_{1}\}$, and $\mathscr{P}_{3}=\{\mathcal{O}_{1},\mathcal{O}_{3}\}$. Set $\mathcal{O}(\mathcal{P}_{i})=\bigcup_{\mathcal{O}_{j}\in\mathscr{P}_{i}}\mathcal{O}_{j}$, and denote $D(\mathscr{P}_{i})=\sum_{\mathcal{O}_{j}\in\mathscr{P}_{i}}|\mathcal{O}_{j}|$ the total number of agents that need to be estimated by local observer on the $i$th agent. Let $Pa(i,j)$ be a set of the nodes belonging to the shortest path between $i$ and $j$ in communication network $\forall i,j\in\mathcal{V}$. For example, in Figure 3c), $Pa(1,4)=\{1,2,3,4\}$. Besides, we define $r_{1},r_{2},\cdots,r_{N}$ for nodes belonging to $\mathcal{V}$, and define the symbol $\preceq$ by $r_{i}\preceq r_{j}$ if $|\mathcal{C}_{r_{i}}|<|\mathcal{C}_{r_{j}}|$, or $|\mathcal{C}_{r_{i}}|=|\mathcal{C}_{r_{j}}|$ and $|\mathcal{N}_{r_{i}}|\geq|\mathcal{N}_{r_{j}}|$. In addition, define $s_{1},s_{2},\cdots,s_{N}$ for nodes belonging to $\mathcal{V}$, and further define $\succeq$ by $s_{i}\succeq s_{j}$ if $|\mathcal{C}_{s_{i}}|>|\mathcal{C}_{s_{j}}|$, or $|\mathcal{C}_{s_{i}}|=|\mathcal{C}_{s_{j}}|$ and $|\mathcal{N}_{s_{i}}|\geq|\mathcal{N}_{s_{j}}|$.
Now, one can see Algorithm 1 for the developed partition method. For the convenience of readers to understand the algorithm, we will briefly describe the algorithm process here.
Step 1: Sort $r_{i},~{}i=1,\ldots,N$ based on symbol $\preceq$.
Step 2: Starting from the minimum $r_{i}$. Each node chooses its partitions according to the greedy algorithm in sequence (The larger the degree of node in the communication network, the more partition schemes can be selected. Therefore, priority should be given to the node with the smallest communication network degree).
Step 3: Sort $s_{i},~{}i=1,\ldots,N$ based on symbol $\succeq$.
Step 4: Starting from the maximum $s_{i}$. Each node considers whether to merge several partitions according to two criteria (Line 27–29 in Algorithm 1 and their details are shown in (D2)) in sequence (In communication networks, nodes with larger degrees are more likely to be forced to join many redundant partitions. Therefore, priority is given to nodes with larger degrees to determine whether to remove some redundant partitions through merging).
Note that the partition algorithms described in Algorithm 1 and Algorithm 2 actually solve a multi-objective optimization problem because each node has its own desired partition result, but the desired optimal partitions among different nodes are conflicting. In the complete version of this article, we prove that the partition results obtained through Algorithm 1 and Algorithm 2 are Pareto solutions to this multi-objective optimization problem.
Propsition 1
Partition results obtained from Algorithm 1 and Algorithm 2 are already located on the Pareto frontier.
Proof:
Given a group of partitions ($\mathcal{O}_{1},\ldots,\mathcal{O}_{m}$) obtained by Algorithm 1 and Algorithm 2. We first prove three propositions.
1) Add any nodes in arbitrary $\mathcal{O}_{i}=\{i_{1},\ldots,i_{r}\},~{}i\in\{1,2,\ldots,m\}$ would increase the dimensions of observers on at least one node. Since $\mathcal{O}_{i}$ is obtained by at least one node (denote as $i_{1}$) within the partition using a greedy algorithm, node $i_{1}$ makes the optimal choice when selecting $\mathcal{O}_{i}$ based on the current situation. Now, let’s discuss two scenarios. First, $i_{1}$ has yet to be included in any partition before selecting $\mathcal{O}_{i}$. In this case, $\mathcal{O}_{i}$ is the first partition for $i_{1}$, so it is the optimal partition choice for $i_{1}$ at that time. In this scenario, adding any new node to this partition will increase the dimension of the observer on $i_{1}$. Second, before selecting $\mathcal{O}_{i}$, $i_{1}$ has already been included in another partition (denote as $\mathcal{O}_{j}$) by another node (let’s call it $i_{2}$). In this case, if a new node $i_{0}$ is added into $\mathcal{O}_{i}$, then the dimension of observers on node $i_{1}$ will increase owing to the joining of $i_{0}$, unless $\mathcal{O}_{j}$ becomes a subset of $\mathcal{O}_{i}\cup\{i_{0}\}$ and node $i$ chooses to merge $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. However, $\mathcal{O}_{j}$ is the optimal choice made by $i_{2}$ based on the greedy algorithm. Therefore, any changes made by $i_{1}$ to $\mathcal{O}_{j}$ will inevitably result in an increase in the observer dimension on $i_{2}$. This proposition is proven.
2) For arbitrary $\mathcal{O}_{i},~{}i\in\{1,2,\ldots,m\}$, every node within this partition is indispensable. Taking $i_{\in}\mathcal{O}_{i}$ as an example, there are two reasons why $i_{1}$ includes $i_{2}$ in the partition based on the greedy algorithm. First, $i_{1}$ needs to estimate the information of $i_{2}$. In this case, if $i_{2}$ is excluded from the partition, $i_{1}$ would still need to establish a new partition $\mathcal{O}_{i}^{\prime}$ that includes both $i_{1}$ and $i_{2}$ in order to estimate the state of $i_{2}$. Thus, it must hold that $|\mathcal{O}_{i}|<|\mathcal{O}_{i}^{\prime}|+|\mathcal{O}_{i}\backslash\{i_{2}\}|=2+|\mathcal{O}_{i}|-1$. Second, if $i_{1}$ does not need to estimate the state of $i_{2}$, it implies that the connectivity of the communication network within $\mathcal{O}_{i}$ depends on the presence of $i_{2}$. In this case, if $i_{2}$ is removed, the partition becomes invalid. In conclusion, we have demonstrated that “any node in any partition is indispensable”.
Based on the above three propositions, we will prove that the partitioning results obtained by Algorithm 1 and Algorithm 2 are already Pareto optimal solutions. To this end, it is sufficient to show that reducing the value of $D(\mathscr{P}_{j})$ for any node (e.g., $j$) will result in an increase in $D(\mathscr{P}_{k})$ for at least one node (e.g., $k$). We will discuss this with two cases.
First, let’s assume that $\mathscr{P}_{j}$ contains multiple partitions ($\mathcal{O}_{j1}$, $\mathcal{O}_{j2}$, $\ldots$). In order to reduce $D(\mathscr{P}_{j})$, node $j$ can either remove a node from a certain partition (e.g., deleting node $k$ from partition $\mathcal{O}_{j1}$) or merge multiple partitions (e.g., merging $\mathcal{O}_{j1}$ and $\mathcal{O}_{j2}$). Based on the previous propositions, there exists at least one node $i_{1}$, for which $\mathcal{O}_{j1}$ is optimal. Therefore, any operation of reducing or merging $\mathcal{O}_{j1}$ will inevitably lead to an increase in $D(\mathscr{P}_{i_{1}})$. Hence, in this case, optimizing the observer dimensions of node $j$ will inevitably result in at least one other node suffering a loss in benefit or resulting in an invalid partitioning result.
Second, let’s assume that $\mathscr{P}_{j}$ contains only one partition. In this case, the only way for node $j$ to reduce $D(\mathscr{P}_{j})$ is by removing some nodes from that partition. However, according to Proposition 2), this approach will obviously result in at least one other node suffering a loss in benefit.
In conclusion, the partition results obtained from Algorithm 1 and Algorithm 2 are already located on the Pareto frontier. ∎
We have the following discussions regarding Algorithm 1 and 2.

### v4

# High Performance Distributed Control for Large-Scale Linear Systems: A Cover-Based Distributed Observer Approach

@@ Abstract
In recent years, the distributed-observer-based distributed control law has shown powerful ability to arbitrarily approximate the centralized control performance. However, the traditional distributed observer requires each local observer to reconstruct the state information of the whole system, which is unrealistic for large-scale scenarios. To fill this gap, This paper presents a coverage solution algorithm for large-scale systems that accounts for both physical and communication network characteristics, which can significantly reduce the dimension of local observers. Then, the cover-based distributed observer for large-scale systems is proposed to overcome the problem that the system dynamics are difficult to estimate due to the coupling between cover sets. Furthermore, the two-layer Lyapunov analysis method is adopted and the dynamic transformation lemma of compact errors is proved, which solves the problem of analyzing stability of the error dynamic of the cover-based distributed observer. Finally, it is proved that the distributed control law based on the cover-based distributed observer can also arbitrarily approximate the control performance of the centralized control law, and the dimension of the local observer is greatly reduced compared with the traditional method. The simulation results show the validity of the developed theories.
@@ I Introduction
A distributed observer comprises a network of $N$ cooperative agents connected via a communication graph. Each agent acts as a node, equipped with a local observer and a local controller. While each local observer only accesses partial system outputs, they collaborate through the communication network to collectively estimate the entire system’s state. Due to their broad applicability—demonstrated in fields such as flexible structures [1], smart vehicles [2], microgrids [3, 4], and spacecraft [5]—and significant theoretical value, distributed observers have attracted considerable research interest [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Theoretically, they enable distributed control laws to achieve performance levels that can arbitrarily approximate centralized control within a sufficiently small error [2, 20, 21, 22, 23], a key advancement in the field.
However, current distributed observer designs suffer from a critical limitation: they require each agent to reconstruct the full state vector of the entire system. This requirement, known as state omniscience, implies that if the global system dimension is $n$, every local observer must also have dimension $n$. For large-scale systems, this becomes computationally prohibitive, contradicting the fundamental goal of distributed design. Although some studies have explored minimum-order [24] or reduced-order observers [25], they typically replace $N$ full-order observers with $N$ lower-order observers. Since the dimension $p_{i}$ of the measurement output of each local observer is much smaller than that of the large-scale system state, the distributed reduced-order observers does not fundamentally solve the scalability issue caused by state omniscience (the dimension of each local observer is $n-p_{i}$).
This paper investigates whether the dimension of local observers can be significantly reduced. The central challenge is to achieve this reduction without compromising the ability of the resulting distributed control law to approach centralized performance. Existing work [26] suggests that the performance gap between distributed and centralized control stems from the former’s limited access to state information. Conversely, distributed observers bridge this gap precisely through state omniscience [2], indicating that state omniscience is a key factor in enabling distributed-observer-based distributed control laws to approximate centralized performance. Therefore, replicating this performance without state omniscience presents a major challenge.
Our approach is inspired by insights from the structure of large-scale interconnected systems. We observe that for such systems—often represented by a physical network graph—the local control law on an agent typically does not require the full states of the entire system. Theoretical analyses demonstrate that by truncating feedback connections beyond a certain topological distance in the physical network, the performance degradation relative to the centralized optimal control is exponentially small. Therefore, the states of its physical network neighbors are sufficient to achieve performance that closely approximates the centralized benchmark. This insight leads to our core question: Can we design a novel distributed observer where each local observer only estimates the states of its physical neighbors? If feasible, the local observer dimension would be drastically reduced, especially since large-scale systems are often sparsely connected. Moreover, if an agent can estimate all its physical neighbors’ states, its local controller can utilize similar information as a centralized one, preserving performance while reducing observer dimension.
Realizing this novel observer is highly non-trivial. A local observer can estimate its physical neighbors’ states directly only if the physical and communication networks are identical—a condition rarely met in practice (For example, smart grid systems, urban water supply systems, etc., their physical networks need to reach thousands of households, but the communication networks responsible for data transmission and control signal issuance can be more flexible and sparse). This mismatch means that a single agent’s communication neighbors may not cover its physical neighbors. Consequently, this paper develops a node covering method where each cover set contains multiple nodes, and nodes can belong to multiple cover sets. A distributed observer is then designed within each cover set. The key is to ensure that for each agent, the union of the cover sets it belongs to covers all its physical neighbors (a necessary condition for approximating centralized performance), while minimizing the total number of involved nodes to reduce local observer dimension.
This leads to the first scientific challenge: designing an intelligent cover solving strategy that meets the above conditions. The literature on coverage or similar partitioning primarily addresses schemes that either divide networks into non-overlapping sections based on specific tasks or requirements [27, 28], optimize partitions against defined objectives [29, 30, 31, 32], or coordinate agents for complete areal coverage [33, 34]. While the third category aligns with our goal, existing methods presuppose known agent parameters and areal coverage, unlike our problem of covering network nodes. Furthermore, in our scenario, both the regions to be covered and the number of covering sets are unknown and must be co-determined through an analysis of the physical and communication networks. Consequently, this divergence renders the literature inadequate for our needs, thus requiring a new algorithmic approach.
The second challenge involves managing the coupled error dynamics across cover sets. Although observers are designed within cover set, their dynamics are coupled with observers in other cover sets. This coupling must be carefully handled in both the observer design and stability analysis. If this coupling relationship is not eliminated, then a distributed observer within the cover sets cannot be designed. Whereas, if the coupling is directly eliminated, then the adverse impacts caused by issues such as model mismatch (there are differences between the model used by the observer and the actual model of the system) must be considered. rendering classical distributed observer theories inadequate for this new architecture.
To address these challenges, this paper makes three key contributions: First, we propose a novel and feasible coverage solving method for large-scale systems, which yields a Pareto-optimal solution for the multi-objective optimization problem inherent in each agent’s conflicting preferences for cover set selection. Second, we present a design method for the cover-based distributed observer and develop a two-layer Lyapunov approach to analyze its error dynamics’ stability. Finally, we demonstrate that the distributed control law based on our cover-based distributed observer can arbitrarily approximate centralized control performance within a sufficiently small error bound, with the performance gap proven to be exponentially small relative to the truncation distance in the physical network topology, providing a theoretical foundation for applying distributed observers to large-scale systems.
@@
The rest of this paper is organized as follows. Section II formulates the problem. Section III details the coverage solving method and observer design. Section IV analyzes the error dynamics and closed-loop performance. Simulation results are presented in Section V, and Section VI concludes the paper.
Notations: Let $\mathbb{R}^{n}$ be the Euclidean space of $n$-dimensional column vectors and $\mathbb{R}^{m\times n}$ be the Euclidean space of $m\times n$-dimensional matrices. Denote $A^{T}$ the transpose of matrix $A$. $\bar{\sigma}(A)$ and $\underline{\sigma}(A)$ stand for the maximum and minimum eigenvalues of $A$, respectively, if $A=A^{T}$. We cast $col\{A_{1},\ldots,A_{n}\}$ as $A_{1}^{T},\ldots,A_{n}^{T}]^{T}$, and $diag\{A_{1},\ldots,A_{n}\}$ as a block diagonal matrix with $A_{i}$ on its diagonal, where $A_{1},\ldots,A_{n}$ are matrices with arbitrary dimensions. $dim\{\cdot\}$ represents the dimension of a vector. $I_{n}\in\mathbb{R}^{n\times n}$ is an identity matrix and $sym\{A\}=A+A^{T}$ if $A$ is a square matrix. We denote $|\cdot|$ the cardinality of a set. $\|\cdot\|$ represents the $2$-norm of vectors or matrices. Furthermore, there are many symbols in this paper. To avoid confusion for readers, we have listed a comparison table of easily confused symbols in Table 1, and corresponding symbols will also be explained when they first appear.
@@
$\mathcal{O}_{p}$ The $p$th cover set
$\mathscr{P}_{i}$ Set of cover sets that containing node $i$
@@
$D(\mathscr{P}_{i})$ $\color[rgb]{0,0,1}\sum_{\mathcal{O}_{j}\in\mathscr{P}_{i}}|\mathcal{O}_{j}|$}
$\hat{x}_{li}^{(p)}$ Estimation of $x_{i}$ generated by observers on agent $l$ in cover set $\mathcal{O}_{p}$
@@ II-A Overall objectives
In traditional distributed observer, each local observer—the orange block in Figure 1—is required to reconstruct all the states of the entire system based on $y_{i}$ and the information exchanged via communication network. Then, each local controller—blue block in Figure 1—can use all the information of the entire system. Therefore, all local observers can jointly establish a distributed control law with centralized performance [2, 21]. This paper will investigate whether it is possible to achieve distributed control laws with arbitrary approximation of centralized control performance without requiring each local observer to estimate global information.
@@
1) Design the cover-based distributed observer with $\hat{x}_{i\star}$ being the states of the $i$th local observer, and $\hat{x}_{i\star}$ satisfies $dim\{\hat{x}_{i\star}\}\ll dim\{x\}$;
@@ II-B An illustrative example for the target problem
Figure 2: Illustrations of coverage issue in simple network systems.
@@
[…]
Up to now, we have briefly described the main idea of this paper through an example. To achieve this idea, there are mainly three steps involved: 1) Design a reasonable and effective coverage solving algorithm; 2) Design cover-based distributed observer in each cover set; 3) Design distributed control law.
@@
The reader should note that Assumption 1 is not completely the same as the traditional assumption in distributed observers. In existing literature [9, 21], it is generally assumed that $(C,A)$ is observable, but $(C_{i},A)$ may not be observable. Assumption 1 used in this paper is given owing to the form of system (1)–(2). This system format is easier to express our core views on covering and status information retrieval. Actually, the cover-based distributed observer can also be achieved based on the General system $\dot{x}=Ax,~y=Cx$ and the traditional assumption as in [9, 21]. However, due to limitations in space and the need to express the main idea, the theory of cover-based distributed observers for general systems will be included in our future research.
III Design of cover-based distributed observer
@@
A large-scale network coverage solving algorithm will be given in this section. Then, based on the coverage solving results, we propose the design method and observer structure of the cover-based distributed observer.
III-A Cover sets selecting
@@
As mentioned in Section I and Subsection II-B, it is unnecessary to require each local observer to reconstruct the states of the whole large-scale system. Theoretically, each agent only needs to estimate part of states required by its local control law. Therefore, in order to establish a distributed observer that can only estimate the state information required by the local control law, we need to find a cover of all agents and ensure that all cover sets of each agent contain all its physical network neighbors (note that in many large-scale systems such as power grids and water networks, communication networks are often distinct from physical networks. This is because communication infrastructure can be flexibly arranged to reduce costs, rather than being constrained by the complex pipeline layouts of physical systems). To illustrate it in more detail, we introduce several simple examples.
Figure 3: Illustrations of covering in simple network systems.
@@
In Figure 3a), there is a simple network system with $6$ nodes. One may recognize its cover by observation easily: all nodes could be covered by three sets including $\{1,2,3\},\{4,5,6\}$, and $\{3,4\}$. With this cover sets, the physical network neighbors of all nodes have been included in their cover sets. For example, $\mathcal{N}_{3}=\{1,2,4\}$, and cover sets to which node $3$ belonging are $\{1,2,3\}$ and $\{3,4\}$. Obviously, $\mathcal{N}_{3}\subset\{1,2,3\}\cup\{3,4\}$. Similarly, Figure 3b) is a star network system with $9$ nodes. One can find a cover as $\{1,2\}$, $\{1,3\}$, $\{1,4\}$, $\{1,5\}$, $\{1,6\}$, $\{1,7\}$, $\{1,8\}$, $\{1,9\}$ and all agents’ physical neighbors are covered by their cover sets. As for Figure 3c), the acquisition of cover sets is not so intuitive, but we can still obtain it with $\{1,2,3,4\}$, $\{5,6,7,8\}$, and $\{3,5\}$ after the observation and analysis. Through the above examples, we understand the basic idea and function of covering. However, for large-scale systems, we cannot obtain a reasonable coverage through mere observation. Therefore, a principled covering algorithm is necessary for distributed observers of large-scale systems.
@@
To this end, we propose a coverage solving method based on the greedy algorithm. Before showing the algorithm details, some notations are introduced. $\mathcal{O}_{k}\subset\mathcal{V}$ is defined as the $k$th cover set of the large-scale system, and $\mathscr{P}_{i}$ denotes the set of cover sets that containing node $i$. For example, in Figure 3c), $\mathcal{O}_{1}=\{1,2,3,4\}$, $\mathcal{O}_{2}=\{5,6,7,8\}$, and $\mathcal{O}_{3}=\{3,5\}$. In addition, $\mathscr{P}_{1}=\{\mathcal{O}_{1}\}$, and $\mathscr{P}_{3}=\{\mathcal{O}_{1},\mathcal{O}_{3}\}$. Set $\mathcal{O}(\mathcal{P}_{i})=\bigcup_{\mathcal{O}_{j}\in\mathscr{P}_{i}}\mathcal{O}_{j}$, and denote $D(\mathscr{P}_{i})=\sum_{\mathcal{O}_{j}\in\mathscr{P}_{i}}|\mathcal{O}_{j}|$ the total number of agents that need to be estimated by local observer on the $i$th agent. Let $Pa(i,j)$ be a set of the nodes belonging to the shortest path between $i$ and $j$ in communication network $\forall i,j\in\mathcal{V}$. For example, in Figure 3c), $Pa(1,4)=\{1,2,3,4\}$. Additionally, we define for node $i$ indices $r_{i}$ and $s_{i}$, with the partial orders $r_{i}\preceq r_{j}$ if $|\mathcal{C}_{i}|<|\mathcal{C}_{j}|$ or ($|\mathcal{C}_{i}|=|\mathcal{C}_{j}|$ and $|\mathcal{N}_{i}|\geq|\mathcal{N}_{j}|$), and $s_{i}\succeq s_{j}$ if $|\mathcal{C}_{i}|>|\mathcal{C}_{j}|$ or ($|\mathcal{C}_{i}|=|\mathcal{C}_{j}|$ and $|\mathcal{N}_{i}|\geq|\mathcal{N}_{j}|$).